Based on the following equation:
6HCL(aq)+ 2AL(s)> 2ALCL3(aq) + 3H2 (g)

What mass of aluminum is needed to produce 500 mL of hydrogen gas at 25 degrees Celcius and 100 kPa?
1
Expert's answer
2020-07-07T14:21:51-0400

P = 100 kPa

T = 25oC + 273 = 298 K

V = 500 mL = 0.5 L

R = 8.314 kPa L K-1 mol-1


Solution:

The balanced chemical equation:

6HCl(aq) + 2Al(s) = 2AlCl3(aq) + 3H2(g)

According to the chemical equation: n(Al)/2 = n(H2)/3


The ideal gas law can be used to calculate the amount of H2:

PV = nRT

n= PV/RT

Thus:

n(H2) = (100 kPa × 0.5 L) / (8.314 kPa L K-1 mol-1 × 298 K) = 0.02 mol


n(Al)/2 = n(H2)/3

n(Al) = 2 × n(H2) / 3 = 2 × (0.02 mol) / 3 = 0.0133 mol

Moles of Al = Mass of Al / Molar mass of Al

The molar mass of Al is 27 g/mol

Thus:

Mass of Al = n(Al) × M(Al) = (0.0133 mol) × (27 g/mol) = 0.3591 g = 0.36 g


Answer: 0.36 g of aluminum (Al) is needed.